1.)The number of positive integer solutions of ab-24=2a is.                                        

Grade:12

2 Answers

arun
7 years ago
given ab-24=2a then, ab-2a=24 this implies a(b-2)=24 now, a=24/(b-2) as (a) is an integer then 24 is divisible by (b-2) and 24=2^3 * 3^1 then total umber of factors are (3+1)(1+1) = 8 so their will be 8 integer value satisfying this equation.
manmath
7 years ago
given ab-24=2a then, ab-2a=24 this implies a(b-2)=24 now, a=24/(b-2) as (a) is an integer then 24 is divisible by (b-2) and 24=2^3 * 3^1 then total umber of factors are (3+1)(1+1) = 8 so their will be 8 integer value satisfying this equation.